OpenStudy (anonymous):

Solve for x. 5x^2 − 80 = 0 {16} {4} {4, −4} {16, −16}

OpenStudy (fibonaccichick666):

so what is the first thing we can do to isolate our x?

OpenStudy (jiteshmeghwal9):

\[5x^2=80\]\[x^2=\frac{80}{5}\]\[x=\sqrt{\frac{80}{5}}=?\]

OpenStudy (jiteshmeghwal9):

solve it further :)

OpenStudy (cggurumanjunath):